OpenStudy (anonymous):

If h(x) = −4x − 10, find h(−5).

OpenStudy (anonymous):

h(-5)=-4(-5) - 10

OpenStudy (anonymous):

h(-5) = 20 - 10; h(-5) = 10

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

10

OpenStudy (anonymous):

just substitute (-5) instead of x in the equation. then h(-5) = 20-10=10